OpenStudy (anonymous):

** I will give you a medal ** Simplify the complex fraction . [(2)/(5t) - (3)/3t)]/[(1)/(2t) + (1)/(2t)]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the LCD of the following fractions 2/(5t) 3/(3t) 1/(2t) ?

OpenStudy (anonymous):

i have no idea...

jimthompson5910 (jim_thompson5910):

what is the LCM of 5t, 3t and 2t

OpenStudy (anonymous):

t ?

jimthompson5910 (jim_thompson5910):

t will be in it, but there's more to it

jimthompson5910 (jim_thompson5910):

i want the LCM, not GCF

OpenStudy (anonymous):

i don't understand..

jimthompson5910 (jim_thompson5910):

the lowest common multiple of 5, 3, and 2 is 30 since 30 is the smallest number that is a multiple of all 3 numbers: 5, 3, 2

jimthompson5910 (jim_thompson5910):

so the LCM of 5t, 3t, and 2t is 30t

OpenStudy (anonymous):

oh okay. i see it.

jimthompson5910 (jim_thompson5910):

therefore, the LCD of these fractions 2/(5t) 3/(3t) 1/(2t) is 30t

jimthompson5910 (jim_thompson5910):

turns out that if you multiply every term and fraction by the LCD, you'll clear out the fractions

OpenStudy (anonymous):

okayy.

jimthompson5910 (jim_thompson5910):

example: 2/(5t) times 30t = 60t/5t = 12 notice how the fraction is now gone

OpenStudy (anonymous):

yeah.. so i do that to all 3 of those fractions?

jimthompson5910 (jim_thompson5910):

yes you do it to every fraction and term

jimthompson5910 (jim_thompson5910):

this is so you balance everything out and don't change anything

OpenStudy (anonymous):

okay so i got 12, 30, and 15

OpenStudy (anonymous):

what do i do now ?

jimthompson5910 (jim_thompson5910):

-(3)/3t times 30t = -90t/3t = -30

jimthompson5910 (jim_thompson5910):

I think that's what you meant

OpenStudy (anonymous):

yes, i forgot the negative sign

jimthompson5910 (jim_thompson5910):

so [(2)/(5t) - (3)/3t)]/[(1)/(2t) + (1)/(2t)] turns into [12 - 30]/[15 + 15]

OpenStudy (anonymous):

so i got -18/30 which equals -9/15

OpenStudy (anonymous):

no.. it equals.. -6/10

jimthompson5910 (jim_thompson5910):

and that can be further reduced

OpenStudy (anonymous):

which equals ... -3/5

jimthompson5910 (jim_thompson5910):

even further

jimthompson5910 (jim_thompson5910):

good, -3/5 is your final answer

OpenStudy (anonymous):

thank you (: (:

jimthompson5910 (jim_thompson5910):

sure thing